OpenStudy (anonymous):

Convert 2.5 g/mL to the unit g/L.

OpenStudy (anonymous):

Hint: \[ \large \left( 2.5\ \frac{\text{g}}{\text{mL}} \right) \underbrace{\left( \color{red}{??}\ \frac{\text{mL}}{\text{L}} \right)}_\text{set to equal 1} =\ \color{red}{??}\ \frac{\text{g}}{\text L} \]

OpenStudy (anonymous):

\[\frac{ 2.5g }{ mL } = \frac{ 2.5g }{ 10^{-3}L }\]

OpenStudy (anonymous):

=??